OpenStudy (anonymous):

Write the equation of a circle with a center at (-5, -1) and a radius of 5.

OpenStudy (anonymous):

(x – 5)2 + (y – 1)2 = 25 (x + 5)2 + (y + 1)2 = 25 (x + 5)2 + (y + 1)2 = 5 (x – 5)2 + (y – 1)2 = 5

OpenStudy (anonymous):

Do you know how to write standard equation of circle?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

is it B?

OpenStudy (anonymous):

It's right! Do you want to know why it's correct?

OpenStudy (asnaseer):

this question is virtually identical in nature to your previous question: http://openstudy.com/updates/4fe0ed47e4b06e92b8701c5d did you understand how the answer was arrived at there?

OpenStudy (anonymous):

@coolaidd We'd like to help you learn the concept one time for all!